i need to see the work step by step for this problem (substitution method):
x=6+y
3x=-y+6

Answer by stanbon(75887)

(substitution method):
x=6+y
3x=-y+6
-----
Substitute for "x" and solve for "y":
3(6+y) = -y+6
18 + 3y = -y+6
4y = -12
y = -3
-----
solve for "x"
x = 6 + y
x = 6 + -3
x = 3
------
Cheers,
Stan H.
==================